Circular No. 8978

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 S1 = 2008 JK (CATALINA-McNAUGHT) B. G. Marsden has identified astrometry of the apparently asteroidal

object 2008 JK (MPS 249095; discovery by A. Boattini on Catalina Sky Survey images, with the first observation tabulated below), which had briefly been posted on the Minor Planet Center’s ‘NEOCP’ webpage last May, as being identical with comet C/2008 S1 (cf. IAUC 8977).

2008 UT  α_{2000}  δ_{2000}  Mag.
May 2.34377  14^{h}55^{m}09.^{s}84  +11^{o}09′28′′.2  19.8

The following orbital elements by Marsden (from MPEC 2008-S26) indicate a relatively close approach to Jupiter (0.18 AU) in Aug. 1990, with the perihelion distance prior to that encounter being q = 1.5 AU.

Epoch = 2008 Oct. 21.0 TT

T = 2008 Oct. 1.8082 TT  ω = 203.^{o}6048
e = 0.667506  Ω = 111.3691 }2000.0
q = 1.190656 AU  i = 15.1108

a = 3.580983 AU n^{o} = 0.1454456 P = 6.78 years

COMET 205P/GIACOBINI D. T. Durig and K. N. Hatchett, Cordell-Lorenz Observatory, Sewa-

nee, TN, have reported observations of two additional components of comet 205P. The first of these, 4 mag fainter than the primary body, was observed on Sept. 17.1, 19.1, and 22.1 UT some 70′′ west and 5′′ south of the primary. The second additional component, 5 mag fainter than the primary body, was observed on Sept. 19.1 and 22.1 UT some 660′′ west and 60′′ south of the primary. Animations of the objects can be seen at

http://arthur.sewanee.edu/obsv/view.php. Approximate ephemerides

for the additional components can be computed by applying corrections of ΔT = +0.014 and +0.133 day, respectively, to the orbital elements for comet 205P on MPC 63822 (or the 2008 set on MPEC 2008-R78, where the provisional designation P/2008 R6 is used).

COMET 29P/SCHWASSMANN-WACHMANN J. A. Henriquez, Tenerife, Canary Islands, reports an outburst of this

comet from CCD images taken with a 0.2-m f/9 reflector; available magnitudes: Sept. 15.18 UT, 14.8; 21.2, 11.3 (stellar appearance).

2008 September 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green